### v2

@@ 3.5 Summary
We outlined the applications of EmAI in healthcare across four pivotal domains: clinical intervention, daily care & companionship, infrastructure support, and biomedical research. These domains are inherently interdependent, collectively forming the cornerstone of human healthcare needs, and EmAI begins to play increasingly vital roles in some of them. As advancements in medical technology and life sciences progress, the boundaries between these domains are becoming increasingly porous. An EmAI system for daily care & companionship may act as an extension of an EmAI system for clinical intervention, and an EmAI with monitoring information for daily care & companionship can offer valuable pre-intervention data for clinical intervention.
In the early stages of EmAI development, limitations in data acquisition and capabilities of multimodal large-scale models, as well as the uneven maturation of EmAI-related research subfields, led to fragmented applications (as we introduced in this section), where an EmAI system could only benefit a specific scenario or sub-scenario. For instance, certain EmAI systems might have been capable of performing image analysis but were not able to plan surgical procedures based on the analysis. However, they share similar backend AI approaches such as world models, LLMs and MLLMs. As EmAI continues to evolve, there is promising potential for these applications to merge, fostering the development of more holistic and versatile EmAI systems capable of meeting the comprehensive needs across all four healthcare domains.
@@ 5 Datasets and Benchmarks
High-quality datasets and benchmarks are indispensable for driving advancements in EmAI research within healthcare. In this section, we highlight representative datasets that are currently used—or hold promise—for training and evaluating EmAI systems in healthcare applications, as illustrated in Figure 11. These datasets span a wide range of application domains, from general clinical procedures to specialized healthcare services, as detailed in Section 3.
For preoperative diagnosis, EmAI models are often trained on diverse medical imaging datasets [721, 722, 723, 724, 725, 726, 727], enabling disease recognition and diagnostic support. In intraoperative procedures, datasets comprising surgical videos or image demonstrations [728, 729, 730, 731, 732, 733, 734] and VLM datasets [735, 736, 737] facilitate multi-modal policy learning and enhance surgical workflows. Additionally, robotic activity and kinematic datasets [738, 739] contribute to advancing surgical automation. As for postoperative rehabilitation, human activity recognition datasets collected via wearable sensors [740, 741, 742, 743] enable automatic healthcare monitoring. Multimodal datasets [744, 745, 746] that integrate vision and sensory inputs empower EmAI systems to analyze patient movements, supporting personalized rehabilitation planning and improving patient outcomes. A critical consideration in curating these datasets is the meticulous collection and de-identification of patient-related data to uphold ethical standards and safeguard privacy [747, 748].
@@
While the existing high-quality datasets have significantly propelled EmAI research in healthcare, several notable limitations hinder their full potential. Surgical procedure datasets like Cholec80 and CaDIS are constrained by relatively insufficient sample sizes, typically comprising a few hundred videos, which restricts the ability to train robust models capable of handling the variability of surgical skills  [728, 730]. Additionally, VLM datasets such as AVOS and LMOD may suffer from inconsistencies in data annotation and synchronization between visual and linguistic inputs, affecting the reliability of multi-modal policy learning [735, 736]. Moreover, Postoperative rehabilitation datasets such as Smart-Insole and UCI-HAPT, while rich in sensor data, often lack comprehensive contextual information necessary for developing personalized rehabilitation protocols [757, 759]. Additionally, biomedical research datasets such as BioRED and Reactome, while comprehensive in biological and chemical information, often face challenges related to data interoperability and standardization, impeding seamless integration with EmAI systems [749, 752]. Datasets that capture instrumental and experimental scenes may not fully reflect the variability of laboratory environments, thereby limiting the adaptability of automated experimental systems [755, 756]. Additionally, the perspective of healthcare world models is highly anticipated, but healthcare data is clearly more limited compared to that in general domains to build a comprehensive world model.
@@ 6.1.1 Liability and Accountability
One of the foremost legal challenges with the implementation of EmAI in healthcare is determining liability in cases of malfunctions or errors that result in patient harm. The question of who is responsible—the device manufacturer, the software developer, or the healthcare provider—becomes complex with AI’s autonomous capabilities [782, 783]. Establishing clear frameworks for liability and accountability is crucial. This involves crafting legislation and regulations that address the unique nature of AI, possibly creating new legal categories that can handle the nuances of AI-driven outcomes [784, 782]. Additionally, insurance policies and liability clauses need to be reevaluated and possibly redesigned to accommodate the role of AI in healthcare, ensuring that patients remain protected and that there is a clear accountability path when things go wrong.
@@ 6.1.2 Ethical Oversight
The deployment of EmAI also raises profound ethical questions, particularly regarding patient consent and the transparency of AI operations [747, 748, 785]. Ethical oversight committees are essential in monitoring and guiding the ethical deployment of AI technologies. These committees should include not only ethicists and legal experts but also technologists, healthcare providers, and patient representatives, to ensure a comprehensive evaluation of AI technologies from multiple perspectives. Key considerations include ensuring informed consent is obtained in a way that patients understand the role and function of AI in their care, and maintaining transparency about how AI uses patient data and how decisions are influenced or made by AI systems [786, 787]. Additionally, ethical oversight must ensure that these technologies do not exacerbate existing healthcare disparities or introduce new forms of bias or discrimination.
@@ 6.2.1 Intervention Security Risks
Beyond the extensively discussed topics of AI security [788, 789, 790], EmAI in healthcare presents an unique challenge – the intervention security risk. The integration of physical devices, such as surgical tools controlled by EmAI-driven decisions, into real-world healthcare scenarios introduces risks of patient harm and medical malpractice. Cyber-attacks [791, 792, 793], adversarial attacks on AI algorithms [794, 795, 796], and data-poisoning attacks [797] can lead to unpredictable outcomes, with direct physical consequences that jeopardize patient safety. Robust protection against unauthorized access and malicious interventions is therefore critical. Moreover, even without external threats, AI algorithms, knowledge databases, and robotic devices within EmAI systems must work in harmony to ensure precise and accurate actions, minimizing errors and safeguarding patient outcomes. This necessitates rigorous validation, real-time monitoring, anomaly detection, and fail-safe mechanisms. While LLM agents acting as evaluators can, to some extent, enhance precision, their performance remains inadequate [798, 799, 800]. Lastly, comprehensive research on the scope of EmAI system application is essential to define security boundaries and ensure appropriate deployment in healthcare settings [90], preventing unintended consequences and ensuring consistent performance—a relatively underexplored area. These challenges underscore the growing need for research in AI security and the precision control of medical devices.
@@ 6.2.2 Data Protection
Another primary concern in healthcare EmAI is ensuring the integrity and confidentiality of data, particularly as EmAI systems may collect comprehensive patient and medical environment data. Encryption techniques are pivotal in this regard. Advanced methods, such as homomorphic encryption, enable AI systems to process encrypted data without the need for decryption, thus preserving privacy even during analysis [801]. Additionally, blockchain technology can be leveraged to create secure, immutable records of medical data transactions, bolstering security across distributed networks [802]. Model pruning techniques [803, 804] that reduce model size also hold potential for enabling offline EmAI systems. These technologies help safeguard against unauthorized access and data breaches, which are critical in patient data management.
@@ 6.2.3 Compliance Systems
Compliance with legal and ethical standards is essential in the deployment of AI in healthcare. Developing compliance systems that incorporate AI can help healthcare providers adhere to complex regulatory frameworks such as the GDPR in Europe or HIPAA in the United States [805, 806]. These systems need to be dynamic, capable of adapting to new regulations and policies as they arise. AI-driven compliance systems can automatically update and audit data use across platforms, ensuring continuous adherence to data protection laws [807]. This not only helps in managing data responsibly but also in building trust with patients and regulatory bodies.
@@ 6.3.1 Accuracy and Explainability Enhancement
AI models, particularly those involved in diagnostics and patient monitoring, must perform with high precision to be trusted by healthcare providers and patients alike. Techniques such as deep learning have shown potential in improving the accuracy of AI systems, but they require large volumes of high-quality, diverse data [808]. Enhancing the reliability of these systems also involves implementing robust validation frameworks that simulate a wide range of clinical scenarios to test AI responses and avoid LLM hallucinations [809, 810]. Continuous learning algorithms can be utilized to update and refine AI models based on new data and outcomes, thus gradually improving their accuracy and reliability over time.
Explainability in EmAI, particularly within the healthcare sector, is crucial for clarifying how these systems perceive, make decisions, and take actions, thereby enhancing both transparency and trustworthiness. Taking a surgical EmAI system as an example, it is essential to present patient condition data and explain each surgical step, such as adjusting a scalpel angle due to abnormal blood flow detection. Such explanations are vital for building clinicians’ trust, especially in the early stages of EmAI adoption when its acceptance is still limited. Therefore, multimodal attribution analysis [811, 812], feature visualization [813], multimodal knowledge graphs [347, 814], causal inference [815, 816, 817], and leveraging the conversational capabilities of LLMs further enhance explainability by providing grounded, theory-based explanations and revealing causal relationships in medical decision-making, thereby aiding clinicians in navigating complex medical scenarios.
@@ 6.3.2 Interoperability of Devices
Interoperability between different AI-powered devices and existing healthcare systems presents another substantial challenge. For EmAI to be effective, it must seamlessly integrate and communicate with various healthcare IT systems and medical devices [818, 819]. Standardization of data formats and communication protocols is crucial to achieve this. Initiatives like Health Level Seven International (HL7) and Fast Healthcare Interoperability Resources (FHIR) standards are pivotal in promoting this integration [820, 821]. By ensuring that AI systems can interact with one another and with legacy systems, healthcare providers can deliver more coordinated and efficient care. Moreover, the development of APIs and open standards can facilitate easier data exchange and functionality across different platforms and devices, enhancing the scalability and utility of AI solutions in healthcare environments. Additionally, the need to develop systems that are compatible with different types of robots and to compress diverse knowledge into one machine brain (e.g., abilities to segment CT, segment X-ray, dialogue capabilities) emphasizes the importance of designing AI with modular architectures that can integrate multiple functionalities and process heterogeneous data efficiently.
@@ 6.3.4 Stronger Generalization Capabilities
The future of EmAI in healthcare lies in its potential to seamlessly integrate physical and AI capabilities, transforming the way healthcare systems interact with patients and their environments. Currently, most research in this field has focused on discrete applications—such as robotic-assisted surgery or rehabilitation—yet the broader frontier is in developing EmAI for more generalized settings. To this end, the exploration of spatial intelligence, multi-modality foundational models, large language models, world models, and physical AI is crucial [352]. Additionally, evolutionary algorithms that autonomously tune hyperparameters and support continuous learning also help models to evolve and better adapt post-deployment in diverse environments.
@@ 6.4.1 User Training and Education
The introduction of EmAI into healthcare settings necessitates significant user training and education to ensure that healthcare providers can effectively interact with and leverage these technologies [833]. Training programs must be developed to cater to a diverse range of skills and familiarity with technology among healthcare workers. These programs should focus on the operational aspects of the technology, as well as on understanding the AI’s decision-making process, to build trust and confidence in AI systems [834]. Furthermore, continuous education needs to be provided to keep pace with the rapid evolution of AI technologies, ensuring that clinicians remain competent in utilizing these advanced tools.
@@ 6.4.2 Workflow Integration
Integrating EmAI systems into existing healthcare workflows presents its own set of challenges. These systems must be designed to complement and enhance existing healthcare practices without causing disruption [835, 836]. For instance, EmAI tools such as robotic assistants, smart prosthetics, and augmented reality systems for surgery or patient care need to be seamlessly integrated into hospital routines. This integration requires careful planning and adjustment of clinical pathways, involving both the redesign of physical spaces and the modification of procedural protocols to accommodate the new technologies. The goal is to ensure that these AI systems improve efficiency and patient outcomes without adding unnecessary complexity to the clinicians’ workload.
@@ 6.5.1 Cost-Benefit Analysis
Implementing EmAI technologies in underdeveloped areas involves significant costs, including the initial investment in technology, ongoing maintenance, and training of clinicians to use these systems effectively [837]. However, the benefits can be substantial, potentially transforming healthcare delivery by providing sophisticated diagnostic tools, telemedicine capabilities, and robotic assistance in regions where clinicians are scarce. A detailed cost-benefit analysis must consider the long-term savings in healthcare costs due to improved disease prevention, diagnostic accuracy, and treatment outcomes. This analysis should also evaluate how these technologies can extend healthcare reach, reduce the burden on understaffed clinics, and improve patient care in remote areas. The ultimate goal is to determine whether the high initial costs are justified by the anticipated improvements in healthcare quality and accessibility.
@@ 6.5.2 Social Resource Allocation
The introduction of EmAI in underdeveloped regions also poses questions about the allocation of social resources. Decisions need to be made about how to distribute limited resources—whether to invest in advanced AI technologies or to allocate funds to more basic healthcare needs like vaccinations and clean water. This becomes a complex issue of prioritization, where the potential long-term benefits of AI must be weighed against immediate healthcare necessities. It is crucial to engage community stakeholders in these decisions to ensure that the allocation of resources aligns with the actual needs and priorities of the community. Moreover, policymakers must consider how these investments in AI technology might affect social equity, ensuring that these innovations do not widen health disparities but instead contribute to more equitable healthcare outcomes [838].

### v3

@@ 3.5 Summary
We outlined the applications of EmAI in healthcare across four key domains: clinical intervention, daily care & companionship, infrastructure support, and biomedical research. These domains are inherently interdependent, collectively forming the cornerstone of human healthcare needs, and EmAI begins to play increasingly vital roles in some of them. As advancements in medical technology and life sciences progress, the boundaries between these domains are becoming more porous. An EmAI system for daily care & companionship may act as an extension of an EmAI system for clinical intervention, and an EmAI with monitoring information for daily care & companionship can offer valuable pre-intervention data for clinical intervention.
In the early stages of EmAI development, limitations in data acquisition and capabilities of multimodal large-scale models, as well as the uneven maturation of EmAI-related research subfields, led to fragmented applications (as we introduced in this section), where an EmAI system could only benefit a specific scenario or sub-scenario. For instance, certain EmAI systems might have been capable of performing image analysis yet could not plan surgical procedures based on the analysis. However, they share similar backend AI approaches such as world models, LLMs and MLLMs. As EmAI continues to evolve, there is promising potential for these applications to merge, paving the way for more holistic and versatile EmAI systems that meet the comprehensive needs across all four healthcare domains.
@@ 5 Datasets and Benchmarks
High-quality datasets and benchmarks are essential for driving advancements in EmAI research within healthcare. In this section, we highlight representative datasets that are currently used—or hold promise—for training and evaluating EmAI systems in healthcare applications, as illustrated in Figure 11. These datasets span a wide range of application domains, from general clinical procedures to specialized healthcare services, as detailed in Section 3.
For preoperative diagnosis, EmAI models are often trained on diverse medical imaging datasets [721, 722, 723, 724, 725, 726, 727], which enable disease recognition and diagnostic support. In intraoperative procedures, datasets comprising surgical videos or image demonstrations [728, 729, 730, 731, 732, 733, 734] and VLM datasets [735, 736, 737] facilitate multi-modal policy learning and enhance surgical workflows. Additionally, robotic activity and kinematic datasets [738, 739] contribute to advancing surgical automation. As for postoperative rehabilitation, human activity recognition datasets collected via wearable sensors [740, 741, 742, 743] enable automated healthcare monitoring. Multimodal datasets [744, 745, 746] that integrate vision and sensory inputs empower EmAI systems to analyze patient movements, supporting personalized rehabilitation planning and improving patient outcomes. A critical consideration in curating these datasets is the meticulous collection and de-identification of patient-related data to uphold ethical standards and safeguard privacy [747, 748].
@@
While the existing high-quality datasets have significantly propelled EmAI research in healthcare, several notable limitations hinder their full potential. Surgical procedure datasets like Cholec80 and CaDIS are constrained by relatively insufficient sample sizes, typically comprising a few hundred videos, which restricts the ability to train robust models capable of handling the variability of surgical skills  [728, 730]. Additionally, VLM datasets such as AVOS and LMOD may suffer from inconsistencies in data annotation and synchronization between visual and linguistic inputs, affecting the reliability of multi-modal policy learning [735, 736]. Moreover, Postoperative rehabilitation datasets such as Smart-Insole and UCI-HAPT, while rich in sensor data, often lack comprehensive contextual information necessary for developing personalized rehabilitation protocols [757, 759]. Additionally, biomedical research datasets such as BioRED and Reactome, while comprehensive in biological and chemical information, often face challenges related to data interoperability and standardization, impeding seamless integration with EmAI systems [749, 752]. Datasets that capture instrumental and experimental scenes may not fully reflect the variability of laboratory environments, thereby limiting the adaptability of automated experimental systems [755, 756]. Additionally, the perspective of healthcare world models is highly anticipated, but healthcare data is clearly more limited compared to that in general domains for building a comprehensive world model.
@@ 6.1.1 Liability and Accountability
One of the main legal challenges with the implementation of EmAI in healthcare is determining liability when malfunctions or errors cause harm to patients. The question of who is responsible—the device manufacturer, the software developer, or the healthcare provider—becomes more complex with AI’s autonomous capabilities [782, 783]. It is essential to establish clear guidelines for liability and accountability. This requires creating laws and regulations that account for the unique nature of AI, possibly introducing new legal categories to address the complexities of AI-driven outcomes [784, 782]. Additionally, insurance policies and liability clauses must be reexamined and potentially redesigned to reflect AI’s role in healthcare, ensuring patient protection and providing a clear path for accountability when things go wrong.
@@ 6.1.2 Ethical Oversight
The deployment of EmAI also raises important ethical questions, particularly around patient consent and the transparency of AI operations [747, 748, 785]. Ethical oversight committees are crucial for monitoring and guiding the responsible use of AI technologies. These committees should include not only ethicists and legal experts, but also technologists, healthcare providers, and patient representatives, to ensure a well-rounded evaluation of AI technologies from various perspectives. Key considerations include making sure that patients fully understand the role and function of AI in their care when giving consent, and being transparent about how AI uses patient data and influences decisions [786, 787]. Ethical oversight must also ensure that these technologies do not worsen existing healthcare disparities or introduce new forms of bias or discrimination.
@@ 6.2.1 Intervention Security Risks
Beyond the extensively discussed topics of AI security [788, 789, 790], EmAI in healthcare presents a unique challenge – the intervention security risk. The integration of physical devices, such as surgical tools controlled by EmAI-driven decisions, into real-world healthcare scenarios introduces risks of patient harm and medical malpractice. Cyber-attacks [791, 792, 793], adversarial attacks on AI algorithms [794, 795, 796], and data-poisoning attacks [797] can lead to unpredictable outcomes, which could result in physical harm to patients and threaten their safety. Robust protection against unauthorized access and malicious interventions is therefore critical. Moreover, even without external threats, AI algorithms, knowledge databases, and robotic devices within EmAI systems must work in harmony to ensure precise and accurate actions, minimizing errors and safeguarding patient outcomes. This necessitates thorough validation, real-time monitoring, anomaly detection, and fail-safe mechanisms. While LLM agents acting as evaluators can, to some extent, enhance precision, their performance remains inadequate for full reliability [798, 799, 800]. Lastly, comprehensive research on the scope of EmAI system application is essential to define security boundaries and ensure appropriate deployment in healthcare settings [90], preventing unintended consequences and ensuring consistent performance—a relatively underexplored area. These challenges underscore the growing need for research in AI security and the precision control of medical devices.
@@ 6.2.2 Data Protection
Another primary concern in healthcare EmAI is ensuring the integrity and confidentiality of data, especially as EmAI systems may collect extensive patient and medical environment data. Protecting this data through encryption is essential. Advanced methods, such as homomorphic encryption, allow AI systems to process encrypted data without needing to decrypt it, thus maintaining privacy even during analysis [801]. Additionally, blockchain technology can be used to create secure, unchangeable records of medical data transactions, further enhancing security across distributed networks [802]. Model pruning techniques [803, 804], which reduce model size, also hold potential for enabling offline EmAI systems. These technologies help protect against unauthorized access and data breaches, which are critical in patient data management.
@@ 6.2.3 Compliance Systems
Compliance with legal and ethical standards is crucial when deploying AI in healthcare. Developing AI-powered compliance systems can help healthcare providers navigate complex regulatory frameworks like the GDPR in Europe or HIPAA in the United States [805, 806]. These systems must be flexible, enabling them to adapt to new regulations and policies as they emerge. By leveraging AI, compliance systems can automatically update and audit data usage across platforms, ensuring ongoing adherence to data protection laws [807]. This not only supports responsible data management but also fosters trust with patients and regulatory bodies.
@@ 6.3.1 Accuracy and Explainability Enhancement
AI models, particularly those involved in diagnostics and patient monitoring, must perform with high precision to be trusted by healthcare providers and patients alike. Techniques such as deep learning have shown potential in improving the accuracy of AI systems, but they require vast amounts of high-quality, diverse data [808]. Enhancing the reliability of these systems also involves implementing robust validation frameworks that simulate a wide range of clinical scenarios to test AI responses and avoid LLM hallucinations [809, 810]. Continuous learning algorithms can be utilized to update and refine AI models as new data and outcomes become available, gradually improving their accuracy and reliability over time.
Explainability in EmAI is particularly critical in healthcare, as it clarifies how these systems perceive, make decisions, and take action—boosting both transparency and trust. For example, a surgical EmAI system should clearly present patient condition data and explain each surgical step, such as adjusting a scalpel angle due to abnormal blood flow detection. These detailed explanations are essential for building trust among clinicians, especially during the early stages of EmAI adoption when acceptance is still growing. To enhance explainability, methods such as multimodal attribution analysis [811, 812], feature visualization [813], multimodal knowledge graphs [347, 814], and causal inference [815, 816, 817] can be employed. Additionally, leveraging the conversational capabilities of LLMs can provide grounded, theory-based explanations, revealing causal relationships in medical decision-making and helping clinicians navigate complex medical scenarios.
@@ 6.3.2 Interoperability of Devices
Ensuring that different AI-powered devices and existing healthcare systems work together smoothly is a major challenge. For EmAI to be effective, it needs to integrate and communicate seamlessly with various healthcare IT systems and medical devices [818, 819]. Standardizing data formats and communication protocols is key to making this possible. Initiatives like Health Level Seven International (HL7) and Fast Healthcare Interoperability Resources (FHIR) standards are crucial in driving this integration [820, 821]. By enabling AI systems to interact with each other and with existing systems, healthcare providers can offer more coordinated and efficient care. Additionally, developing APIs and open standards can simplify data exchange and functionality across different platforms and devices, making AI solutions more scalable and useful in healthcare settings. It’s also important to create systems that work with different types of robots and can combine various capabilities—such as segmenting CT scans, X-rays, or engaging in dialogue—into a single system. This suggests the need for AI with modular designs that can integrate multiple functions and handle diverse data efficiently.
@@ 6.3.4 Stronger Generalization Capabilities
The future of EmAI in healthcare lies in its potential to seamlessly integrate physical and AI capabilities, transforming the way healthcare systems interact with patients and their environments. Currently, most research in this field has focused on discrete applications—such as robotic-assisted surgery or rehabilitation—yet the broader frontier is in developing EmAI for more generalized settings. To this end, the exploration of spatial intelligence, multi-modality foundational models, large language models, world models, and physical AI is crucial [352]. Additionally, evolutionary algorithms that autonomously tune hyperparameters and support continuous learning also help models evolve and better adapt post-deployment in diverse environments. By enabling a network of autonomous agents to coordinate and share real-time insights, EmAI systems can further develop more adaptive and scalable generalization capabilities through the power of collective intelligence.
@@ 6.4.1 User Training and Education
The introduction of EmAI into healthcare settings requires significant user training and education to ensure that healthcare providers can effectively interact with and leverage these technologies [833]. Training programs must be developed to cater to a diverse range of skills and familiarity with technology among healthcare workers. These programs should focus on the operational aspects of the technology, as well as on understanding the AI’s decision-making process, to build trust and confidence in AI systems [834]. Furthermore, continuous education needs to be provided to keep pace with the rapid evolution of AI technologies, ensuring that clinicians remain competent in using these advanced tools.
@@ 6.4.2 Workflow Integration
Integrating EmAI systems into healthcare workflows is challenging but necessary. These systems must be designed to complement and enhance existing healthcare practices without causing disruption [835, 836]. For instance, EmAI tools such as robotic assistants, smart prosthetics, and augmented reality systems for surgery or patient care need to be seamlessly integrated into hospital routines. This integration requires careful planning and adjustment of clinical pathways, involving both the redesign of physical spaces and the modification of procedural protocols to accommodate the new technologies. The goal is to ensure that these AI systems improve efficiency and patient outcomes without adding unnecessary complexity to the clinicians’ workload.
@@ 6.5.1 Cost-Benefit Analysis
Deploying EmAI technologies in underdeveloped areas comes with significant costs, including the initial investment in technology, ongoing maintenance, and training of clinicians to use these systems effectively [837]. However, the benefits can be substantial, as EmAI can revolutionize healthcare delivery by offering advanced diagnostic tools, telemedicine, and robotic assistance in areas with a shortage of clinicians. A detailed cost-benefit analysis must consider the long-term savings in healthcare costs due to improved disease prevention, diagnostic accuracy, and treatment outcomes. This analysis should also evaluate how these technologies can extend healthcare reach, reduce the burden on understaffed clinics, and improve patient care in remote areas. The ultimate goal is to determine whether the high initial costs are justified by the anticipated improvements in healthcare quality and accessibility.
@@ 6.5.2 Social Resource Allocation
The introduction of EmAI in underdeveloped regions also poses questions about the allocation of social resources. Decisions need to be made about how to distribute limited resources—whether to invest in advanced AI technologies or to allocate funds to more basic healthcare needs like vaccinations and clean water. This becomes a complex issue of prioritization, where the potential long-term benefits of AI must be weighed against immediate healthcare necessities. It is crucial to engage community stakeholders in these decisions to ensure that the allocation of resources aligns with the actual needs and priorities of the community. Moreover, policymakers must consider how these investments in AI technology might affect social equity, ensuring these technologies reduce health disparities rather than widening them [838].